OpenStudy (anonymous):

NEED HELP RIGHT AWAY PLEASE I WILL GIVE A MEDAL TOO WHO HELPS 3*4+3*3*8 A.312 or B.168

Mehek (mehek14):

\(3*4=12\\3*3=9*8=72\\12+72=84\) how come that's not a choice?

OpenStudy (anonymous):

i am not sure cause that is the same thing i got

Mehek (mehek14):

yea it has to be 84 if you follow the order of operations

OpenStudy (anonymous):

maybe this may help place grouping symbols in the expression to get expressions that have the given values.

OpenStudy (anonymous):

Can you please post a photo of the question. Possibly a transcription error?

OpenStudy (anonymous):

\[3*4+3^{2}*8\]

OpenStudy (anonymous):

there you go just figured out how to do that sorry

OpenStudy (anonymous):

So using order of operations, start with the 3^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then the two sets of multiplication

OpenStudy (anonymous):

Then go for the addition last

Mehek (mehek14):

\(3*4+3^2*8\\3^2=9\\3*4=12\\12+9*8\\9*8=72+12=84\) same thing